Baka sold article for $6900 and made a profit of 15%. Calculate his profit if he had sold it for $ 6600
Answer by Edwin McCravy(20055)

There are two quantities that somebody could mean by the word "profit".
There's money profit and there's percentage profit.

X = how much Baka paid for the article.

0.15 = his percentage profit expressed as a decimal

0.15X = his money profit

Add the money he paid and his profit and get how much he sold it for:

    X + 0.15X = 6900
1.00X + 0.15X = 6900
        1.15X = 6900
            X = 6900/1.15
            X = 6000      <--what Baka paid for the article

If he had sold it for only $6600, the money value of his profit would have been
$6600 - $6000 = $600.

To find what percentage $600 is of $6000 we divide 600/6000 = 0.10 = 10%.

The money profit would have been $600 and the percentage profit would have been
10%.

Edwin